# The Impact of Surface Co-location and Eye-tracking on Mixed Reality Typing

### v2

@@ 3.2. Procedure
Participants used a HoloLens 2 headset while seated at a small table (Figure 1 left). We told participants they could type using the index fingers on one or both hands. During the calibration phase, participants used only one hand to avoid potential depth measurement issues caused by sensing the non-typing hand. During the practice and evaluation sentences, we reminded users they could type with both hands. We allowed participants the freedom to type with one or both hands as it better simulates real-world keyboard typing where users have the same freedom.
@@
The study lasted approximately one hour. Participants completed an initial questionnaire before starting the conditions. Between conditions, participants took a two-minute break. During the break, we instructed participants to lift the HoloLens visor to reduce eyestrain. They then completed a questionnaire about the previous condition. After all conditions, participants completed a final questionnaire. On the questionnaires, participants ranked various statements on a 7-point Likert scale (1=strongly disagree, 7=strongly agree). We also asked for open responses to various questions. For our exact questionnaires, see our supplementary materials.111https://osf.io/q4n72 Participants received a $15 USD gift card.
@@
The order of conditions was completely counterbalanced between participants. Prior to each condition, participants typed a calibration sentence to ensure the keyboard’s consistent placement relative to the surface. If the keyboard was too far in front of the surface, the finger could actuate a key without tactile feedback. If it was too close to the surface, the tracked finger would hit the real surface before it actuated a virtual key—making it impossible to type. Calibration computed the average depth of nine key presses when the participants typed “calibrate”. The participants were instructed to keep pushing the key down until they felt the surface. While the calibration sentence only affected the Wall and Table conditions, we retained it in the Midair condition for consistency.
@@
Following calibration, participants typed two practice sentences and twelve evaluation sentences. We only analyzed the evaluation sentences. Sentences were pulled randomly without replacement from the “mem1-5” set of the mobile Enron dataset (Vertanen and Kristensson, 2011). These sentences were in mixed case and included punctuation and sometimes numbers.
Eighteen participants between the ages of 18 and 26 took part in Experiment 1. 15 participants identified as male, 2 as female, and 1 as non-binary. 16 participants strongly agreed with the statement “I consider myself a fluent speaker of English”, one participant agreed, and one participant strongly disagreed. We believe the participant who strongly disagreed had made an error when marking the scale. Eight participants reported augmented reality (AR) experience, including one participant who reported owning a HoloLens. The remaining seven participants reported limited AR experience, primarily with phone apps such as Pokémon Go.
@@
Throughout this paper, we test for statistical differences using repeated measures ANOVA. In cases were Mauchly’s test indicated a violation of sphericity, we report Greenhouse-Geisser corrected tests. For significant omnibus tests, we test for differences between conditions via Bonferroni corrected post-hoc tests. For Likert ratings, we used Friedman’s ANOVA. For significant omnibus tests, we test for differences between conditions via a Bonferroni corrected Wilcoxon signed-rank test.
@@ 3.3. Results
Entry rate was measured in words-per-minute (WPM). We considered a word to be five characters (including space). We subtracted one from the number of characters to compensate for the unmeasured time finding and typing the first key (Wobbrock, 2007). We started timing from when the first key was pressed. We ended timing when the next button was pressed since this includes the perceptual, cognitive, and motor overheads associated with confirming the last character typed. Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively (Figure 2 left). This difference was significant (Table 1 left). Post-hoc tests found Midair was significantly faster than both Table and Wall.
@@
Pairwise | Midair Table ( ¡ ) | n/a | n/a |
post-hoc | Midair Wall ( ¡ ) | ||
Table Wall ( = ) |
We measured typing errors using character error rate (CER). CER refers to the number of character insertions, substitutions, and deletions required to change a participant’s final entered text to the reference text, relative to the characters in the reference and multiplied by 100. Error rates were low in all conditions with a CER of 0.95%, 0.98%, and 0.65% in Midair, Table, and Wall respectively (Figure 2 center). This difference was not significant (Table 1 center).
As a measure of how much effort participants were expending correcting errors, we calculated the number of backspaces per final output character for each phrase. While the final character error rates in the conditions were relatively low, we observed frequent use of backspace. The number of backspaces pressed per output character was 0.12 for Midair, 0.16 for Table, and 0.13 for Wall. While there did seem to be more backspacing in the two surface co-located keyboards (in particular for Table), this difference was not significant (Table 1 right).
@@
We recorded the position of participants’ left and right index fingers 30 times per second. We also noted whenever tracking was lost for either finger. During the time periods participants were typing the evaluation phrases, we computed the percentage of a participant’s samples in which at least one finger was tracked. We found tracking success was 99.9%, 99.1%, and 98.6% in Midair, Table, and Wall respectively. While this difference was significant ( = , , ), post-hoc pairwise comparisons were not significant: Midair Table ( = ), Midair Wall ( = ), and Table Wall ( = ). We note that just because we had a tracked location of the index fingers, this does not mean that location was necessarily accurate. Subjective feedback from many participants indicated hand-tracking was less accurate in Midair and Table. We conjecture this could have resulted from a number of factors such as the proximity of the hands to the surface, occlusion of the fingers or hands between themselves, or changes due to the relative location and angle between the hands and the headset.
@@
To investigate if participants’ accuracy was impacted by the spatial location of their desired key, we calculated the percentage of time participants successfully obtained the next letter in the target phrase. We calculated this only on key presses in which, at the time of the key press, a participant’s current typing result was a strict prefix of the target. For example, if the target phrase was “Are you there?” and the participant had typed “Are you t” we would score the next key press based on a target of “h”. However if they had typed “Are you h” we would drop the next key press as we are unsure if they accidentally missed the “t” key or if they are writing an incorrect word like “here”. Filtering in this way allows us to be reasonably confident about a participant’s next intended key. As shown in Figure 4, there did not seem to be a strong difference in accuracy based on spatial location in Midair or Table. However in the Wall condition, the bottom keyboard row was noticeably less accurate at 85% versus 91% for the middle row and 92% for the top row.
@@ 3.4. Subjective feedback
After each condition, participants rated three statements on a 7-point Likert scale (Figure 5). Participants’ average ratings for “I entered text quickly” were: Table 4.0, Wall 2.9, and Midair 5.1. The difference between ratings was significant (). The only significant post-hoc test was that Midair was quicker than Wall (). Participants’ average ratings for “I entered text accurately” were: Table 4.5, Wall 3.8, and Midair 5.4. The difference between ratings was significant (). The only significant post-hoc test was that Midair was more accurate than Wall (). Participants’ average ratings for “Entering text was physically easy and not tiring” were: Table 4.3 for, Wall 2.9, and Midair 4.4. The difference between ratings was significant (). The only significant post-hoc test was that Midair was physically easier than Wall ().
After each condition, we asked participants what they most liked and most disliked about the condition. Table 2 shows selected comments about each condition. Overall, a common like in Table and Wall was the tactile feedback afforded by the physical surface. However, in both conditions, a common dislike was problems related to the tracking of hands and failure to register key presses. Participants complained of physical strain in all conditions, though Table received the most positive feedback regarding ergonomics. For all participant comments, see our supplementary materials.222https://osf.io/q4n72
Condition | Comment | |
Midair | + | “Easier to lean arms against table.” |
+ | “Not needing to touch the surface behind me helped me to lift my fingers up farther, reducing misinput.” | |
+ | “My hands wanted to be there naturally and didn’t require as much movement between letters.” | |
- | “Because my hands were in the air, it made physically more exhausting for longer period of time.” | |
- | “There’s no tactile feedback.” | |
- | “Without the wall or table it was difficult to estimate the depth of the keyboard.” | |
Table | + | “I got to keep my arms on the table.” |
+ | “It was a more natural typing posture.” | |
+ | “Having the table under the keyboard allowed for a form of tactile response.” | |
- | “It seemed I lost tracking a lot in this part; the 3D outlines of the keys felt disorienting.” | |
- | “Accurate input required me to fully focus on my hands, causing back strain.” | |
- | “I had to look further down than normal.” | |
Wall | + | “I liked that the keyboard would type pretty much right as I hit the wall.” |
+ | “Using this interface, I tended to forget that the keyboard wasn’t physically in front of me.” | |
+ | “I didn’t have to pull away so much between letters. I didn’t really touch the table but touched the wall more.” | |
- | “Awkward position, tiring.” | |
- | “Hard to visualize where I needed to put my hands to type best. Typing felt very finicky.” | |
- | “I felt like the keyboard strokes were more inaccurate than the other two tests.” |
@@
In a final questionnaire, participants were asked to rank the conditions. Of the 18 participants, 13 preferred Midair, 5 preferred Table, and none preferred Wall. We asked participants to explain their ranking of the conditions. Of those who ranked Midair the best, some representative comments were: “midair was the fastest and tracked most accurately”, and “hovering was most comfortable and easy to use”. Of those ranking Table as the best, participants commented: “horizontal was natural feeling and responsive”, and “horizontal provided the most comfort and accuracy for me”. Most participants ranked Wall as the worst and participants often mentioned physical strain: “vertical was the slowest and most physically taxing”, and “wall was neither comfortable nor accurate”.
@@ 3.5. Discussion
Our goal in Experiment 1 was to assess whether typing with tactile feedback outperformed or was preferred compared to typing in midair. We conjectured that the tactile feedback conditions, especially the Table condition which is similar to a standard physical keyboard, would perform better and be more preferred. However, our data showed the Midair condition had a significantly higher entry rate than both Table and Wall. All conditions had low final error rates, though moderate use of backspace was observed. Furthermore, in the final questionnaire, thirteen of the eighteen participants preferred the Midair condition, with the second most preferred condition being the Table condition by the remaining five people. We observed throughout the study that the Table and Wall conditions reduced the reliability of key actuation and this was corroborated by comments made by participants. This may explain the increased performance and preference for the Midair condition.
@@
We believe the poor performance in the Table condition may be due to our calibration process or the hand-tracking. During calibration, the participant had to type the word “calibrate”. We instructed participants to continue pushing the key until their finger touched the surface. This form of calibration was dependent on how precise the participant was. We found that with quick hand movements, the HoloLens’ hand-tracking lagged and when the hand abruptly stopped, the resulting depth measurement could be deeper than the participant’s finger had actually gone. The latency of the hand-tracking may also have caused inaccuracies, Kim et al. (Kim et al., 2023) report a latency of 90 ms and tracking jitter of 1 mm using a HoloLens 2 headset. It might be possible to improve the keyboard’s placement above the surface with a modified or extended calibration procedure. It might also be advisable to allow users to interactively adjust the keyboard height during typing. A different tracking system may have produced different results. However, this work provides valuable information about how well surface co-location works given the tracking of a commonly available MR headset.
@@
We also speculate the unreliable key triggering is related to limitations of the HoloLens’ egocentric cameras or its hand-tracking models. During the experiment, we saw participants try to compensate for problems with hand-tracking in a number of ways, such as moving closer to the keyboard or adjusting the viewing angle between the headset and the keyboard. However, in some cases, such adjustments only made tracking less reliable. Ideally, the HoloLens interface would provide some sort of informative feedback that could guide users to adjust their position or environment to improve tracking. Presently, the only feedback is the unstable rendering of a user’s hand skeleton.
@@
Midair interaction can be tiring since users have to hold their hands up with little support (Boring et al., 2009; Hincapié-Ramos et al., 2014). Given this, we were surprised to find participants rated Midair the least tiring (Figure 5 bottom). As we expected, the Table was also comfortable, likely because people could rest their arms on the table similar to traditional keyboard typing. The Wall condition was reported to be the most physically demanding. Although participants could rest their elbows on the table in this condition, it might have worked better for some participants than others. This condition is also the most different from traditional typing. For consistency with the other conditions, participants stayed seated at the table while typing on the wall. It could be that wall typing would be more ergonomic while standing.
@@ 4. Experiment 2
Experiment 1 showed a significantly higher entry rate for the midair keyboard and users strongly preferred that keyboard. The majority of participants also requested the ability to type with multiple fingers. Based on this, we decided to focus exclusively on a midair keyboard in Experiment 2. Our goals were to 1) explore if ten-finger typing is feasible on a midair keyboard, and 2) determine if eye-tracking could aid in eliminating accidental key presses.
@@ 4.1. Design
In pilot testing of ten-finger typing, we found users had difficulty maintaining a consistent hand position in midair. Monitoring the positions of all ten fingers was difficult and this resulted in a high number of accidental key presses. The lack of tactile feedback also meant users often did not realize they had unintentionally pressed a key. Because of this, we worried that ten-finger typing, on its own, would not be competitive with index finger typing.
@@
Like Experiment 1, the keyboard was deterministic without auto-correct or word predictions. In the conditions using the eye-tracking filtering, we enabled all keys within a 2.6 cm radius of the user’s gaze location (roughly a one key radius). Keys outside this radius were disabled. Disabled keys were gray, while enabled keys were transparent (Figure 7). When pressed, a key flashed a slightly transparent blue and a sound was played. All keyboard text was white.
@@
A key was triggered when a participant’s finger pushed a key downward past a threshold of approximately one-third of a centimeter. In the event that multiple keys were pressed at the same time, only the furthest pressed key was registered as pressed. Following a key press, there would be a half-second cool down where the keyboard would not register any key presses to help prevent multiple keys from being triggered from a single tap gesture.
We used a QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace. The keyboard was 35 cm 12 cm, with each letter key being 2.2 cm 2.2 cm. Red spherical colliders were attached to the user’s index fingers in Index and IndexEye, and all ten fingers in Ten and TenEye.

### v3

@@ 3.2. Procedure
Participants used a HoloLens 2 headset while seated at a small table (Figure 1 left). The table was 91 cm wide and 51 cm deep and we placed a QR code at its top left corner. At the start of each condition, participants had to look at the QR code. This provided a spatial anchor to allow consistent placement of the virtual keyboard in each condition. We implemented our spatial anchoring using a MR toolkit extension (see our source code for details).
@@
The study lasted approximately one hour. Participants completed an initial questionnaire with demographic questions, questions about their text entry experience, and questions about their experience with VR/AR. Between conditions, participants took a two-minute break. During the break, we instructed participants to lift the HoloLens visor to reduce eyestrain. Participants then completed a questionnaire about the previous condition. Participants ranked on a 7-point Likert scale (1=strongly disagree, 7=strongly agree) whether they felt the condition allowed them to enter text 1) quickly, 2) accurately, and 3) physically easily. We also solicited freeform feedback on what they liked and disliked about the previous condition. After all conditions, participants completed a final questionnaire in which they ranked conditions from best to worst and explained their ranking. We also asked them what they thought needed improving in AR in general. For our exact questionnaires, see our supplementary materials.222https://osf.io/q4n72 Participants received a $15 USD gift card.
@@
The order of conditions was completely counterbalanced between participants. Prior to each condition, participants typed a calibration sentence to set the keyboard’s height above the surface. If the keyboard was too far in front of the surface, the finger could actuate a key without tactile feedback. If it was too close to the surface, the tracked finger would hit the real surface before it actuated a virtual key—making it impossible to type. Calibration computed the average depth of nine key presses when the participants typed “calibrate”. The participants were instructed to keep pushing the key down until they felt the surface. While the calibration sentence only affected the Wall and Table conditions, we retained it in the Midair condition for consistency. During calibration, participants used only one hand to avoid potential depth measurement issues caused by sensing the other hand.
@@
Following calibration, participants typed two practice sentences and twelve evaluation sentences. We only analyzed the evaluation sentences. Sentences were pulled randomly without replacement from the “mem1-5” set of the mobile Enron dataset (Vertanen and Kristensson, 2011). These sentences were in mixed case and included punctuation and sometimes numbers. During the practice and evaluation sentences, we reminded users they could type with both hands.
3.2.1. Independent variables and statistical tests.
@@
Experiment 1 had a single independent variable, namely the location of the keyboard (midair, table, or wall). We tested for statistical differences using a one-way repeated measures ANOVA. In cases were Mauchly’s test indicated a violation of sphericity, we report Greenhouse-Geisser corrected tests. For significant omnibus tests, we test for differences between conditions via Bonferroni corrected post-hoc tests. For Likert ratings, we used a non-parametric Friedman’s ANOVA test. For significant omnibus tests, we test for differences between conditions via a Bonferroni corrected non-parametric Wilcoxon signed-rank test.
3.2.2. Participants and demographics.
We recruited 18 participants aged 18–26. 15 participants identified as male, 2 as female, and 1 as non-binary. 16 participants strongly agreed with the statement “I consider myself a fluent speaker of English”, one participant agreed, and one participant strongly disagreed. We believe the participant who strongly disagreed had made an error when marking the scale. Eight participants reported augmented reality (AR) experience, including one participant who reported owning a HoloLens. The remaining seven participants reported limited AR experience, primarily with phone apps (e.g. Pokémon Go333Pokémon Go is an AR game released in 2016 in which players locate, capture, train, and battle virtual creatures that appear in real-world locations.).
@@ 3.3. Results
3.3.1. Entry rate.
Entry rate was measured in words-per-minute (WPM). As is common in text entry research, we considered a word to be five characters (including space) (Wobbrock, 2007). We subtracted one from the number of characters to compensate for the unmeasured time of finding and typing the first key (Wobbrock, 2007). We started timing from when the first key was pressed. We ended timing when the next button was pressed since this includes the perceptual, cognitive, and motor overheads associated with confirming the last character.
Entry rates were 12.2, 9.1, and 7.8 WPM in the Midair, Table, and Wall conditions respectively (Figure 2 left). This difference was significant (Table 1 left). Post-hoc tests found Midair was significantly faster than both Table and Wall.
@@
Pairwise | Midair Table, ¡ | n/a | n/a |
post-hoc | Midair Wall, ¡ | ||
tests | Table Wall, = |
3.3.2. Error rate.
We measured typing errors using character error rate (CER). CER refers to the number of character insertions, substitutions, and deletions required to change the final entered text to the reference text, relative to the characters in the reference and multiplied by 100. Error rates were low in all conditions with a CER of 0.95%, 0.98%, and 0.65% in Midair, Table, and Wall respectively (Figure 2 center). This difference was not significant (Table 1 center).
3.3.3. Correction rate.
As a measure of how much effort participants were expending correcting errors, we calculated the number of backspaces per final output character for each phrase. While the final character error rates in the conditions were relatively low, we observed frequent use of backspace. The number of backspaces pressed per output character was 0.12 for Midair, 0.16 for Table, and 0.13 for Wall. While there did seem to be more backspacing in the two surface keyboards (in particular for Table), this difference was not significant (Table 1 right).
@@
3.3.4. Hand-tracking success.
We recorded participants’ left and right index finger positions 30 times per second, including noting whether tracking was lost for either finger. We computed the percentage of samples where at least one finger was tracked. Participants’ tracking success percentages during the evaluation phrase were 99.9%, 99.1%, and 98.6% in Midair, Table, and Wall respectively. While this difference was significant ( = , , ), post-hoc pairwise comparisons were not significant: Midair Table ( = ), Midair Wall ( = ), and Table Wall ( = ). We note that just because we tracked both fingers, this does not mean the returned location was necessarily accurate. As we will discuss, subjective feedback from many participants indicated hand-tracking was less accurate in Midair and Table. We conjecture this could have resulted from a number of factors such as the proximity of the hands to the surface, occlusion of the fingers or hands between themselves, or changes due to the relative location and angle between the hands and the headset.
@@
3.3.5. Accuracy over keyboard.
To investigate if accuracy was impacted by the spatial location of a key, we calculated the percentage of time participants successfully obtained the next letter in the target phrase. We calculated this only on key presses in which, at the time of the key press, a participant’s current typing result was a strict prefix of the target. For example, if the target phrase was “Are you there?” and the participant had typed “Are you t” we scored the next key press based on a target of “h”. However if they had typed “Are you h” we would drop the next key press as we are unsure if they missed the “t” key or if they are writing an incorrect word like “here”. This filtering allowed us to be reasonably confident about a participant’s next intended key. As shown in Figure 4, there did not seem to be a strong difference in accuracy based on spatial location in Midair or Table. However in the Wall condition, the bottom keyboard row was noticeably less accurate at 85% versus 91% for the middle row and 92% for the top row.
@@ 3.4. Subjective feedback
After each condition, participants rated three statements on a 7-point Likert scale (Figure 5). Participants’ average ratings for “I entered text quickly” were: Table 4.0, Wall 2.9, and Midair 5.1. The difference was significant (). The only significant post-hoc test was that Midair was quicker than Wall (). Participants’ average ratings for “I entered text accurately” were: Table 4.5, Wall 3.8, and Midair 5.4. The difference between ratings was significant (). The only significant post-hoc test was that Midair was more accurate than Wall (). Participants’ average ratings for “Entering text was physically easy and not tiring” were: Table 4.3 for, Wall 2.9, and Midair 4.4. The difference between ratings was significant (). The only significant post-hoc test was that Midair was physically easier than Wall ().
3.4.2. Condition ranking.
@@
We asked participants to rank the conditions and to explain their reasoning. Of the 18 participants, 13 preferred Midair, 5 preferred Table, and none preferred Wall. We asked participants to explain their ranking of the conditions. Of those who ranked Midair the best, some representative comments were: “midair was the fastest and tracked most accurately”, and “hovering was most comfortable and easy to use”. Of those ranking Table as the best, participants commented: “horizontal was natural feeling and responsive”, and “horizontal provided the most comfort and accuracy for me”. Most participants ranked Wall as the worst and participants often mentioned physical strain: “vertical was the slowest and most physically taxing”, and “wall was neither comfortable nor accurate”.
3.4.3. Likes and dislikes.
We asked participants what they most liked and most disliked about the system after each condition. One author performed qualitative coding on these responses and another peer reviewed the codings and made minor improvements. For all participant comments, see our supplementary materials.444https://osf.io/q4n72
Fatigue was mentioned by 8, 6, and 2 participants in the Wall, Midair, and Table conditions respectively. There was mixed feedback about the tactile feedback. 8 participants said they liked the tactile feedback in the Wall condition and 4 liked it in the Table condition. For example, one participant liked that they could “touch the wall instead of nothing” and another liked the “physical sensation of touching the table.” None complained about the the presence of haptic feedback in those conditions. However, in the Midair condition, three people reported liking no tactile feedback and one disliked not having tactile feedback.
We counted people who mentioned poor tracking in some form (unexpected key presses, failing to detect key presses, or poor tracking in general). 9 people reported poor tracking in Wall, 5 reported it in Midair, and 13 people reported it in Table. Comments disliking tracking in the Table condition included “it seemed I lost tracking a lot in this part” and “sometimes I would hit one key to the left or right of the intended key.”
3.4.4. AR improvements.
@@ 3.5. Discussion
Our goal in Experiment 1 was to assess whether a keyboard positioned on a surface outperformed or was preferred compared to typing in midair. We conjectured that the keyboards located on surfaces, especially the Table condition which is similar to a standard physical keyboard, would perform better and be more preferred. However, our data showed the Midair condition had a significantly higher entry rate than both Table and Wall. All conditions had low final error rates, though moderate use of backspace was observed. Furthermore, 13 of the 18 participants preferred the Midair condition, while the remaining 5 preferred the Table condition. Although Wall was ranked last, it was also the condition with the most people mentioning how they liked the tactile feedback. We observed throughout the study that the Table and Wall conditions reduced the reliability of key actuation and this was corroborated by participant feedback. This may explain the increased performance and preference for the Midair condition.
@@
Poor performance in the Table condition may be due to our calibration process or the hand-tracking. During calibration, the participant typed the word “calibrate”. We instructed participants to continue pushing the key until their finger touched the surface. This form of calibration was dependent on how precise the participant was. We found that with quick hand movements, the HoloLens’ hand-tracking lagged and when the hand abruptly stopped, the resulting depth measurement could be deeper than the participant’s finger had actually gone. The latency of the hand-tracking may also have caused inaccuracies, Kim et al. (2023) report a latency of 90 ms and tracking jitter of 1 mm using a HoloLens 2 headset. It might be possible to improve the keyboard’s placement above the surface with a modified or extended calibration procedure. It might also be advisable to allow users to interactively adjust the keyboard height during typing. A different tracking system may have produced different results. However, this work provides valuable information about how well position a keyboard on a surface works given the tracking of a commodity MR headset.
@@
We also speculate the unreliable key triggering were related to limitations of the HoloLens’ egocentric cameras or its hand-tracking models. During the experiment, we saw participants try to compensate for hand-tracking issues in a number of ways, such as moving closer to the keyboard or adjusting the viewing angle between the headset and the keyboard. However, in some cases, such adjustments only made tracking less reliable. Ideally, the HoloLens interface would provide some sort of informative feedback that could guide users to adjust their position or environment to improve tracking. Presently, the only feedback is the unstable rendering of a user’s hand skeleton.
@@
Midair interaction can be tiring since users have to hold their hands up with little support (Boring et al., 2009; Hincapié-Ramos et al., 2014). While we positioned our midair keyboard such that participants could brace their forearms against the table edge, the like/dislike qualitative feedback suggested that the Midair condition was nearly as fatiguing as the Wall condition. However, we were surprised to find participants rated Midair the least tiring in the Likert responses (Figure 5 bottom). The Likert and like/dislike data both show that Table was the most comfortable, likely because people could rest their arms on the table similar to traditional keyboard typing. Both sets of data also show that the Wall condition was the most physically demanding. Although participants could rest their elbows on the table in this condition, it might have worked better for some participants than others. This condition is also the most different from traditional typing and placed the keyboard the furthest away from the user. For consistency with the other conditions, participants stayed seated at the table while typing on the wall. Wall typing might be more ergonomic while standing.
@@ 4. Experiment 2
Experiment 1 showed a significantly higher entry rate for the midair keyboard and users strongly preferred that keyboard. The majority of participants also requested the ability to type with multiple fingers. In Experiment 2, we focus exclusively on a midair typing. Our goals were to 1) explore if ten-finger typing is feasible on a midair keyboard, and 2) determine if eye-tracking could aid in eliminating accidental key presses.
@@ 4.1. Design
In pilot testing of ten-finger typing, we found users had difficulty maintaining a consistent hand position in midair. Monitoring the positions of all ten fingers was difficult and this resulted in a high number of accidental key presses. The lack of tactile feedback also meant users often did not realize they had unintentionally pressed a key. Because of this, we worried that ten-finger typing, on its own, would not be competitive with index-finger typing.
@@
Like Experiment 1, the keyboard was deterministic without auto-correct or word predictions. In the conditions using the eye-tracking filtering, we enabled all keys within a 2.6 cm radius of the user’s gaze location. This effectively enabled the key the user was looking at plus all adjacent keys (Figure 6). Keys outside this radius were disabled. Disabled keys were gray, while enabled keys were transparent. When pressed, a key flashed a slightly transparent blue and a sound was played. All keyboard text was white.
@@
A key was triggered when a participant’s finger pushed a key downward past a threshold of approximately one-third of a centimeter. In the event that multiple keys were pressed at the same time, only the furthest pressed key was registered as pressed. Following a key press, there was a half-second cooldown where the keyboard would not register any key presses to help prevent multiple keys from being triggered from a single tap gesture.
We used a QWERTY keyboard including numbers, symbols, shift keys, caps lock, and backspace. The keyboard size (including four rows of characters plus spacebar) was 35 cm 12 cm. Each letter key was 2.2 cm 2.2 cm. Red spherical colliders were attached to a user’s index fingers in Index and IndexEye, and all ten fingers in Ten and TenEye.